.A biconvex lens of focal length 20 cm is cut out into two plano-convex lenses. The focal length of each part is a) 10 cm b) 20 cm c) 30 cm d) 40 cm

Grade:upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The focal legth of each part is doubled ie 40cm

Regards
Vikas TU
3 years ago
f = 20 
p = 100/f = 5D 
P1 = P2 = 2.5 D 
Power of combination = 2.5 + 2.5 = 5D 
f = 100/5 = 20cm …...................